OpenStudy (anonymous):

Given the system of constraints, name all vertices of the feasible region. Then find the maximum value of the given objective function. x>=0 y>=0 y<=1/3x+3 5>=y+x Objective Function: C = 6x – 4y

OpenStudy (anonymous):

@sammixboo please help... I've been on this question forever

OpenStudy (anonymous):

@wormie2014 @tester97 @iambatman @oOKawaiiOo @perl Someone please help

OpenStudy (anonymous):

i'm not sure what this answer is

OpenStudy (anonymous):

@wormie2014 Okay, thank you :)

OpenStudy (anonymous):

@StudyGurl14 sorry to bother you when you first get online but can you please help me?

OpenStudy (anonymous):

@StudyGurl14 If I were to plug in points to the objective function.. would it look something like this? C = 6(0,0)(5,0) – 4(0,3) ?

OpenStudy (studygurl14):

Your first step would be to graph all the equations on the same graph and see the region in which the graphs overlap

OpenStudy (anonymous):

Yes I graphed them but I need help after that..... here is the link:

OpenStudy (anonymous):

https://www.desmos.com/calculator/ygtpnkixcp

OpenStudy (studygurl14):

Okay, so take those points, and then plug each of them into the objective function. Which ever one results in the highest amount of C, is the answer

OpenStudy (studygurl14):

Remember, (x,y)

OpenStudy (anonymous):

C = 6(0,0)(5,0)4(5,0) What now? and do I plug in (1.3,3.5)?

OpenStudy (studygurl14):

Plug in one at a time and solve (0,0) You already know this one is going to be the lowest, so don't even bother pluggin this one in (5,0) C = 6x – 4y C = 6(5) - 4(0) See?

OpenStudy (studygurl14):

(refresh if you see question marks)

OpenStudy (anonymous):

Oh okay and calculate each one?

OpenStudy (studygurl14):

yep

OpenStudy (anonymous):

So for the first one... C = 30?

OpenStudy (anonymous):

And do this for all of them including (1.5, 3.5) ?

OpenStudy (studygurl14):

yep

OpenStudy (studygurl14):

correct, for that one, C = 30

OpenStudy (anonymous):

So whichever is the highest or least is the answer?

OpenStudy (studygurl14):

Which ever is the highest is the answer since you are tying to find the maximum of the objective function

OpenStudy (anonymous):

And thats it? :D

OpenStudy (studygurl14):

yep :)

OpenStudy (anonymous):

WOW thank you I have been on this question since yesterday you are so smart I love you!!! :'D

OpenStudy (studygurl14):

lol, thank you. :)

OpenStudy (anonymous):

THANK YOU! <3

OpenStudy (studygurl14):

anytime :)

OpenStudy (anonymous):

For anyone with this problem, answer is 30!! :)